Circular No. 6696

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970616 A. Dey, Kitt Peak National Observatory; and D. Stern and A. Bunker,

University of California at Berkeley, report: “A spectrum of the optical transient identified by Galama et al. (IAUC 6687) in the field of GRB 970616 was obtained using the Low Resolution Imaging Spectrograph (LRIS) on the Keck II Telescope on July 2.58 UT. This reveals the object to be a late-type M dwarf. Comparison with the atlas by Kirkpatrick et al. (1991, Ap.J. Suppl. 77, 417) suggests a spectral type of M8.5V–M9V. No Hα emission is observed. LRIS images obtained on July 1.60 UT in a 2′′.5-aperture diaphragm yield R = 24.5 ± 0.2, I = 21.1 ± 0.1.”

COMET C/1996 B2 (HYAKUTAKE) M. J. Mumma and V. A. Krasnopolsky, NASA Goddard Space Flight

Center; and M. Abbott, Center for Extreme Ultraviolet Astrophysics, Berkeley, report: “We searched for soft x-ray radiation from comet C/1996 B2 on 1996 Mar. 21–24 (r = 1.11–1.05 AU, Δ = 0.15–0.105 AU), using the Extreme Ultraviolet Explorer orbiting observatory. An image of soft xrays in the range 70–180 eV was obtained with the Deep Survey Camera. It shows crescent-shaped emission with total luminosity of 1.0×10^{25} photon/s within an aperture radius of 120 000 km. The emission is offset by 42 000 km sunward from the nucleus in the sky plane. Spectra were acquired over the range 7–70 nm (180–18 eV) and show both spectral lines and continuum emission.”

COMET 81P/WILD 2 Total visual magnitude estimates: Mar. 27.17 UT, 9.8 (C. E. Spratt,

Victoria, BC, 0.20-m reflector); Apr. 3.79, 9.8 (K. Sárneczky, Raktanya, Hungary, 20×60 binoculars); 7.85, 9.7 (A. Kammerer, Ettlingen, Germany, 0.20-m reflector); 10.09, 9.8 (J. Bortle, Stormville, NY, 0.41-m reflector); 13.80, 9.7 (V. Znojil, Brno, Czech Republic, 25×100 binoculars); 21.81, 9.8 (Znojil); 29.90, 10.2 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); May 5.10, 10.2 (Bortle); 10.88, 9.8 (M. Plsek, Lelekovice, Czech Republic, 0.20-m reflector); 25.92, 10.9 (C. Vitorino, Cabo da Roca, Portugal, 0.25m reflector); 31.88, 10.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); June 4.88, 10.4 (M. Lehky, Hradec Kralove, Czech Republic, 0.42m reflector); 10.88, 10.9 (Lehky); 28.86, 11.2 (Znojil); July 1.46, 11.0 (A. Pearce, City Beach, Western Australia, 0.2-m reflector); 3.46, 11.1 (Pearce).

1997 July 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden